Are the graphs of the lines in each pair parallel? Explain.
y = -3/2x + 2
3x + 2y = 8

Answer by lwsshak3(11628)

Are the graphs of the lines in each pair parallel? Explain.
y = -3/2x + 2
3x + 2y = 8
**
Standard form of equation for a straight line: y=mx+b, m=slope, b=y-intercept
Rewrite given equations in standard form:
First equation: y=-3x/2+2
Second equation:
3x + 2y = 8
2y=-3x+8
y=-3x/2+4
..
As you can see, both lines have the same slope but different x-intercepts, therefore, lines are parallel to each other